40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2260     Entry time: Mon Dec 3 10:44:19 2018
Author: anchal 
Type: Notes 
Category: TempCtrl 
Subject: Plan of action for installing new heater shield and temperature sensors 

We have come to the conclusion that to properly tune PID, we need to accurately model a physical system that represents the cavity thermodynamics and use it to tune PID coefficients. For this, we need good temperature sensors on the cavity (not present right now) and better heat actuation with known control. Awade also mentioned that we have new gold plated heat shields ready to be installed. So we have decided to replace the heat shield and install new temperature sensors as well. Following is the plan of action for the same:

  1. Collect/order required components:
    Platinum 1kOhm RTDs (From Cryo lab).
    REF200 for the current source for RTDs (From Cryo Lab)
    Glue for sticking RTDs to heat shield: VME Varnish (Name might be wrong, ask Aaron. From Cryo Lab).
    Appropriate transimpedance amplifier for AD590 (see ATF:2250) in Vacuum Can Temperature sensors.
    Kapton tape heater stickers (PSL lab)
    New Heat Shields (PSL Lab)
    Vacuum compatible wires (PSL Lab)
    D-sub crimp tool (Maybe Koji has it or look for it in other labs)
    New feedthrough cap with more pins (2x2 for heater pads + 4x2 for RTDs) (see this link, check with Aaron and 40m also)
  2. Temperature sensor circuit:
    Design and solder new temperature sensor circuit. Refer to Aaron's designs (sent in email).
    Test the circuit with RTDs.
    Solder about 3 inch long vacuum compatible wires (2 at each end) of RTDs with male D-sub pins crimped to their ends.
  3. Heater pads and heat shield:
    Determine the size and number of heater pads required in each heat shield.
    Using a new blade, scrape off the insulation on top of pads of heater stickers.
    Solder about small length vacuum compatible wires at the pads of heater stickers. If two pads need to be connected, connect them in series using such small wires.
    Crimp male D-sub pins to the ends of the wires.
    Install the pads on the inside of heat shields making sure to avoid the opening for cavity supports.
  4. New Vacuum Can Temperature Sensor circuit:
    Modify the circuit (maybe solder a new one on a prototype board) for AD590 temperature sensors.
    Use ATF:2250 to select a new opamp or look for better ones.
    We want to reduce thermal drifts and keep input-referred current noise of the opamp low.
    Insulate the new box well with thermocol shields and keep connectors identical.
  5. Assemble and testing:
    Using a very small amount of varnish, stick RTDs to the heat shield on the outer side.
    If possible and if feedthrough cap allows, add two back-up RTDs on the heat shield as well.
    Test the temperature sensors and heater pads and record data.
    Using mass and material properties of heat shield (by calculating specific heat), get an estimate of heater pad's heating power as a function of current sent.
    Test the new temperature sensor box for vacuum can by replacing it with the existing circuit.
  6. Replacing old shields and sensors:
    Disconnect all wires going into the vacuum can.
    Clear up space on the table (behind south PMC, there is some space) for temporarily storing optics.
    Take off the beatnote detection setup by lifting the platform out.
    Open the vacuum can (following its own procedures).
    Take out the cavities with the heat shields.
    Stick aluminum tape on the inside of vacuum can. It is currently dark colored inside.
    Replace the heat shields with the new ones.
    With the appropriate feedthrough cap, crimp vacuum compatible wires of good length (to keep slack) to the feedthrough cap.
    Crimp female D-sub pins at the end of these wires and connect them to heater pads and RTDs.
    Place the cavities with shields on back in the can and close the can. Pump down.
  7. Aligning again:
    Put back the beatnote platform back in place.
    Align the laser into the cavities again. (Use PSL:2253 for notes)
    Align the output from cavities to overlap the beams properly.
    Test if we are back to comfortably locking and moving beams in beatnote frequency space.
  8. Plant model and PID tuning:
    Do Step response tests to figure out cavities specific heat and conductivity with Vacuum Can and each other.
    Treating Cavities and Vacuum can as point masses, model a set of differential equations representing thermodynamics of cavities. (See PSL:2191 and PSL:2194)
    Write a code to simulate the model and work as a fake plant for the PIDlocker_beta.py
    Run some tests to see if the fake plant responds similar to the real plant and modify it accordingly.
    Run PIDlocker_beta.py with the fake plant to tune the PID coefficients faster than with actual plant.
    Test with the tuned parameters. Tune. Test. Repeat.

Hopefully, by the end of this work, we get better stability on beatnote frequency locking to get good spectrum reading at higher resolution from Marconi.

ELOG V3.1.3-